OpenStudy (anonymous):

What is the factorization of the trinomial below? 4x2 - 20x + 24

OpenStudy (anonymous):

hmm maybe this time try it without the steps ^^ do you have problems with the method?

OpenStudy (anonymous):

Yes with the A*C because I'm mixed up about which is which

OpenStudy (anonymous):

It is:\[ax^2+bx+c\]

OpenStudy (anonymous):

A*C? you have three ways to solve this: the square method the cross method i showed earlier and the general equation method choose and i'll help^^

OpenStudy (anonymous):

the cross method

OpenStudy (anonymous):

The a-c method: 1. Find factors of a*c that add up to b-term

OpenStudy (anonymous):

sure, hang on a sec just note though, its a good and fast method, but its harder once the numbers get bigger, and some equations won't be able to be solved by this

OpenStudy (anonymous):

then use gcf

OpenStudy (anonymous):

 try it yourself this time, i'll guide you through!

OpenStudy (anonymous):

basically the first column you'll need to find two numbers that when multiplied, get you the thing below, which is 4x^2 so there are two possibilities: x^2 4x^2 and 2x^2 2x^2 this method is actually a trial and error method, so you'll need practice.

OpenStudy (anonymous):

dang i gotta go, message me if anything

OpenStudy (whpalmer4):

\[4x^2-20x+24\]Start by identifying any common factors: \[4*x^2-4*5x+4*6 = 4(x^2-5x+6)\]We'll set that \(4\) aside until later, and just factor the stuff inside the parentheses.\[x^2-5x+6\] To factor \(x^2-5x+6\), we need to find a pair of factors of \(6\) that will add up to \(-5\). Clearly, both will have to be negative to give us a positive \(6\) as the product and \(-5\) as the sum. We can factor 6 into two negative factors as follows: \[-1*-6\]\[ -2*-3\]\[ -3*-2\]\[ -6*-1\] Clearly, our choice is \(-2, -3\), so our factoring would be\[x^2-5x+6 = (x-2)(x-3)\]Combining that with the common factor of \(4\) we found earlier, \[4x^2-20x+24=4(x-2)(x-3)\] Let's check our answer: \[4(x-2)(x-3) = (4x-8)(x-3) = 4x*x - 3*4x -8*x -8(-3) \]\[ =4x^2-12x-8x+24\]\[=4x^2-20x+24\checkmark\]